OpenStudy (anonymous):

Find the solution of (x-1) (x-2) (x-3) = 0

OpenStudy (anonymous):

just put each bracket equal to zero e.g. x-1=0 then x=1

OpenStudy (anonymous):

x-1=0 x-2=0 x-3=0 x-1+1=0+1 x-2+2=0+2 x-3+3=0+3 x=1 x=2 x=3 So the solution set is {1,2,3}

OpenStudy (anonymous):

yes